OpenStudy (anonymous):

Solve each equation by completing the square. If necessary, round to the nearest hundredth.

OpenStudy (anonymous):

33. x^2+10x=-8 35. x^2-14x+16=0

OpenStudy (anonymous):

Solve each equation using any method. Explain why you chose the method you used. 45. x^2-6x+9=0

OpenStudy (mathmale):

First: be certain that you understand what "completing the square" entails. You start with a quadratic expression or equation and re-write it in different form, so that the new form contains the square of a binomial. If, as in #33, you have the quadratic equation \[x ^{2} + 10x=-8,\] your first goal is to rewrite \[x ^{2}+10x\] in the form \[(x+5)^{2}=25-8.\] How? See the coefficient of the x term in the original equation? It's 10. Take half of this coefficient, that is, take half of 10, and square the result.

OpenStudy (mathmale):

Add that result to both sides of the original equation. You'll get: \[x ^{2}+10x+25-25=-8.\] Now the first three terms can be re-written as \[\left( x+5 \right)^{2}.\] Then we'll end up with \[\left( x+5 \right)^{2}\]-25=-8.

OpenStudy (anonymous):

Ok thank you.

OpenStudy (mathmale):

So...all set, as is?

OpenStudy (anonymous):

Huh?

OpenStudy (mathmale):

I'd typed a lot more, but lost it. I wanted to know whether you're now able to finish solving the problem yourself.

OpenStudy (anonymous):

Yes.

OpenStudy (mathmale):

Great. just out of curiosity, where are you located? I'm in Southern California.

OpenStudy (anonymous):

South Carolina

OpenStudy (mathmale):

Cool. Good luck! See you.